OpenStudy (logan13):

What is the value of the expression 3–4? A. –81 B. –12 C. – D.

OpenStudy (logan13):

http://prntscr.com/a6pghx

OpenStudy (logan13):

@alex6799

rebeccaxhawaii (rebeccaxhawaii):

3-4=-1

rebeccaxhawaii (rebeccaxhawaii):

Please click "best response" for the person that helped you the most, this just shows the appreciation for the time they took to assist. If you want further help with anymore questions, please fan me so in the long run you can see when I'm on. I'll be sure to take a look at your question to help :)

OpenStudy (logan13):

alex

OpenStudy (logan13):

there is no negative 1 the link shows the rest of the answers

OpenStudy (alex6799):

\[3^{-4}\] \[\frac{ 1 }{ 3^{4} }\] \[\frac{ 1 }{ 81 }\]

rebeccaxhawaii (rebeccaxhawaii):

you need to make it 3^-4 to show its an exponent

rebeccaxhawaii (rebeccaxhawaii):

but yes alex is correct

OpenStudy (logan13):

so its D

OpenStudy (alex6799):

yes